OpenStudy (stryder12345):

PLEASE HELP WILL GIVE MEDAL Choose the correct simplification of the expression 4 over y to the power of negative 3. A 4y3 B 4 over y to the 3rd power C y to the 3rd power over 4 D This expression is already simplified.

OpenStudy (stryder12345):

Please help

OpenStudy (calculusxy):

\[\large x^{-a} = \frac{1}{x^a}\] or the inverse is \(\huge \frac{1}{x^{-a}}\) = \(\huge x^a\) `4 over y to the power of negative 3` 1) There is a fraction, where 4 is the numerator and \(y^{-3}\) is in the denominator. \(\large \frac{4}{y^{-3}}\). Since a negative power is raised, it will be put to the numerator, where it would become: \(\large \frac{4y^3}{1}\) or \(\large 4y^3\)

OpenStudy (stryder12345):

so it is A

OpenStudy (calculusxy):

Yes